Backhaul Machine movers depreciate their equipment over a 5-year-period using the straight-line depreciation method. On January 1, 2014, the equipment was purchased for $120,000. Its residual value is zero. What will be the net book value of the equipment by the end of December 2016?

Select one:

a. $72,000

b. $96,000

c. Not enough information is provided to calculate a value

d. $48,000
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(120,000/5)=$24,000/year

Hence book value by the end of December 2016=Cost-Accumulated Depreciation

=120,000-(24000*3)

=$48000